OpenStudy (anonymous):

tanx+tany=a, cotx+coty=b Pt cot(x+y) =1/a +1/b?

hartnn (hartnn):

use the cot(A+B) formula : \(\Large \cot(A+B) = \dfrac{\cot A \cot B -1}{\cot A+\cot B}\)

hartnn (hartnn):

you already have the denominator use tan x + tan y = a convert tan to cot

OpenStudy (anonymous):

Will try and ping...

hartnn (hartnn):

although you will get cot(x+y) =1/a -1/b so please verify your question....

OpenStudy (anonymous):

{(1/tanx*1/tany) - 1} /b How to solve?

hartnn (hartnn):

tanx+tany=a 1/cot x+1/cot y = a (cot y +cot x)/ (cot y cot x) = a b/ (cot y cot x) = a got this ?

OpenStudy (anonymous):

Ya...

OpenStudy (anonymous):

Cotycotx is (1/tanx) (1/tany) right?

hartnn (hartnn):

yes but lets keep it as cot y cot x b/ (cot y cot x) = a cot y cot x = b/a right ?

hartnn (hartnn):

now you have numerator too

OpenStudy (anonymous):

Thanks a lot! :) :)

hartnn (hartnn):

welcome ^_^